A) A student measures the molar solubility of calcium carbonate in a water solution to be 6.35×10-5 M.
Based on her data, the solubility product constant for this compound is
B) A student measures the CrO42- concentration in a saturated aqueous solution of calcium chromate to be 2.74×10-2 M.
Based on her data, the solubility product constant for calcium chromate is
Answer #1

A)

At equilibrium:

CaCO3 <----> Ca2+ + CO32-

   s s

Ksp = [Ca2+][CO32-]

Ksp = (s)*(s)

Ksp = 1(s)^2

Ksp = 1(6.35*10^-5)^2

Ksp = 4.032*10^-9

Answer: 4.03*10^-9

B)

At equilibrium:

CaCrO4 <----> Ca2+ + CrO42-

   s s

Ksp = [Ca2+][CrO42-]

Ksp = (s)*(s)

Ksp = 1(s)^2

Ksp = 1(2.74*10^-2)^2

Ksp = 7.508*10^-4

Answer: 7.51*10^-4